What is the probability that 40 or more boxes will be delivered on a particular day?


1
Expert's answer
2022-03-30T13:39:15-0400

from the distribution

Number of Boxes (X) Probability P(X)

35 1/10

37 1/5

40 1/10

42 1/5

45 3/10

50 1/10

the probability that 40 or more boxes will be delivered on a particular day will be calculated as

P(X ≥ 40)

so the value of x are P(X ≥40)= P(40) + P(42) +P(45) +P(50)

P(X ≥ 40) =1/10 +1/5 + 3/10 +1/10= 7/10 or 0.70